# GrAFFORD: A Benchmark Dataset for Testing the Knowledge of Object Affordances of Language and Vision Models

### v1

@@ 4 Experiments
We explore various state-of-the-art baselines using pre-trained language models (RoBERTa-large, BART-large), instruction-fine-tuned large language models (e.g., FLAN-T5, Falcon), pre-trained multi-modal vision and language architectures (CLIP-ViT, ViLT, InstructBLIP, IDEFICS, LLaVA). We observe whether these models gain the knowledge of affordances through their pre-training, fine-tuning on commonsense tasks (NLI, PIQA), or few-shot fine-tuning scenarios.
@@ 4.1.1 Pre-trained language models
Zero-shot prediction task is framed in different ways.
Masked language modelling (MLM) based approach. Here, we pose the zero-shot task as masked word prediction problem. Talmor et al. (2020) performed zero-shot prediction by employing Multi-Choice Masked Language Modelling approach, where the pre-trained model are required to predict the masked word from a set of given keywords. We choose BERT-large-uncased, RoBERTa-large Zhuang et al. (2021), and BART-large Lewis et al. (2020) models for the experiment. We pass the sentence and prompt separated by a [SEP] token as an input to the model. We use the prompt “<Object> can be used for <MASK_TOKEN> by human” and obtain the probability of each affordance label using the logit corresponding to the <MASK_TOKEN>.
Predictions from generative LLMs. Petroni et al. (2019); Schick and Schütze (2021) treats NLU tasks as a cloze test using prompts. We apply autoregressive language models such as FLAN-T5 Chung et al. (2022) (large, xl, and xxl), Falcon Almazrouei et al. (2023) (7b and 40b), ChatGPT to get the predictions. We provide with a ‘YES/NO’ question-answer based prompt to the LLMs to predict whether a particular affordance can be performed on the given object. Based on rigorous prompt engineering we choose specific prompts for the different models as shown in the Appendix Table 7. We map ‘YES/NO’ predictions to 1/0 labels respectively.
@@ 4.1.2 Common sense reasoning tasks
To understand whether the injection of the common sense knowledge in the pre-trained models can enhance the performance of the affordance prediction, we first fine-tune the pre-trained models on common sense reasoning dataset such as PiQA Bisk et al. (2019). Then we run the fine-tuned models on our dataset using the MLM setup. We use BERT-base, BERT-large, RoBERTa-large, and BART-large in this setup.
Natural language inference (NLI) based approach. The NLI task considers a premise and a hypothesis as input pair $\langle p,h\rangle$, and the models are trained to predict the probability whether the hypothesis is entailed by, contradicts or neutral with respect to the premise. Here we use the entailment probability from the models: $p_{La}(h|p)=p(l=``ENTAILMENT"|(p,h)).$ This approach requires language models to be fine-tuned on premise-hypothesis pairs with the corresponding labels. Here we use RoBERTa-large and BART-large fine-tuned on the Multi-genre NLI (MNLI) corpus Williams et al. (2018) consisting of 433k sentence pairs. For each sentence-object pair in our dataset, we use the corresponding context sentence in the MNLI dataset as the premise. We use the hypothesis as “<object> can be used for <affordance> by human” for each object present in the sentence and 15 affordance classes. Using the NLI setting, we predict the entailment score for each affordance class for the given sentence-object pair. We use these scores for ranking the affordance classes and report mAP scores as well as accuracy.
4.1.3 Multi-modal models
@@
Unimodal text-only MLM setup
@@
Images contain necessary information about shape, texture, and size (visual attributes) of objects that can be utilized to effectively predict an object affordance (such as the handle of the bucket can be used to grasp and lift). Hence, we also convert the problem into a multi-modal task by retrieving (or generating) a corresponding image from the context sentence, and predict the affordance of an object (mentioned in the sentence) based on the input.
Retrieving images. In this setup, we use two different techniques to retrieve semantically close images to corresponding context sentences using 1) retrieval and 2) generation. We further use top five images for both, to get an accurate estimation.
Retrieval based: Yang et al. (2022) used Bing Image Search [^6] https://learn.microsoft.com/en-us/azure/cognitive-services/bing-image-search/overview to retrieve images based on object. However, this process is costly as it requires paid subscription. So, we employ Visualgenome Krishna et al. (2017) dataset, consisting of 108,077 images and  3.8 million object instances as the image database. We first encode the images using multi-modal CLIP Radford et al. (2021) based sentence-transformers architecture, and index those image embeddings using Approximate Nearest Neighbour search (ANN) [^7] https://pypi.org/project/annoy/, for making the search efficient. Now, for each sentence, we search for top five images from the database to be used further.
We use the top five retrieved images by using retrieval and generation methods each. We use CLIP Radford et al. (2021) and ViLT Kim et al. (2021) as our vision-text models. CLIP is pre-trained on 400M image-caption pairs with the contrastive learning strategy. CLIP has a text encoder $f_{T}$ and a visual encoder $f_{V}$, which can project text and image into the shared latent space. We aggregate the $k$ (=5) corresponding images and use CLIP to compute the relevance score of (x, y): $Score_{VI}(x,y)=\frac{1}{k}\sum_{i=1}^{K}{\cos{(f_{T}(x),f_{v}(I^{k}_{y}))}}$, where $I^{k}_{y}$ is the $k^{\textrm{th}}$ image for the input text $y$. ViLT uses patch projection Dosovitskiy et al. (2021) to encode images, and uses image-text matching, MLM, and image-patch alignment tasks as objectives. For the zero-shot prediction, we provide the text prompt along with the representative images as input to the ViLT model to predict the masked token. We use the same prompt as the previous MLM task (i.e., “<Object> can be used for <MASK_TOKEN> by human.”) and get the probability of each affordance class as the logit corresponding to the <MASK_TOKEN>.
Text generation based. Similar to section 4.1.1, we utilize state-of-the-art VLMs to make predictions regarding object affordances. We provide with a ‘Yes/No‘ question answering based text prompt along with the aligned images as input to the VLMs, and the model should generate an answer whether a particular affordance can be performed on the given object. We use state-of-the-art VLMs such as IDEFICS Laurençon et al. (2023), LLaVA Liu et al. (2023), InstructBLIP Dai et al. (2023) for this task. The text prompt used for the models can be found in Table 7.
4.2 Ensemble language and vision prediction
@@ 4.3 Few-shot prediction
Training data To create few-shot training examples for fine-tuning encoder based PTLMs, we take all the 62 objects, and for each object we randomly select exactly 1 positive affordance class (i.e., the class label annotated as 1) and 1 negative affordance class (i.e., the class label annotated as 0) for generating the training prompt. As this dataset does not contain any context sentences for a corresponding object, we use ChatGPT UI to generate the sentences for the corresponding objects and manually verify the sentences, so that it does not contain any invalid information. Finally, we have 62 sentence-object pairs and 2 classes (one positive and one negative) per pair, which we use to generate training examples. Each training example consists of a prompt and a label. They constitute 124 training examples (62 sentence-object pairs and 2 selected classes for each) for the few-shot experiment.
Selecting examples for in-context learning: We randomly sample five sentence-object-affordance triples from the above training data as the incontext demonstration examples in such a way that there should be $k$ positive affordance classes. We vary the number of positive affordance classes $k\in\{1,2,3\}$ and report the average accuracy.
@@
5 Result
@@
For each positive class $\in\{P_{1},P_{2},..P_{n}\}$, we compare the predicted scores of that affordance class with the predicted scores of the negative classes $\in\{N_{1},N_{2},..N_{m}\}$. If the predicted score of the positive class is higher than the predicted score of all the negative classes (i.e., $p(P_{i})>p(N_{j})_{\forall j}$), we increment the correct count by 1888During calculation we discard the cases when there is no positive class for a sentence-object pair in the ground truth. We do not find any instance where no negative class is present.. Conversely, if the predicted score of the negative class is higher, we increment the wrong count by 1. The final accuracy is calculated by dividing the total number of correct counts by the total number of the instances. To rank the affordance classes based on the predicted score, we also report the Mean Average Precision (mAP@$K$, where $K$ is the number of affordance classes).
@@
NLI based (Normal)
Actual LM + VI (CLIP) LM + VI (ViLT)
Model Acc mAP Acc mAP Acc mAP
RoBERTa-large-mnli 0.64 0.43 0.79 0.52 0.79 0.54
BART-large-mnli 0.65 0.38 0.62 0.4 0.64 0.43
@@
Model Acc mAP Acc mAP Acc mAP
BERT-large-uncased 0.46 0.26 0.55 0.38 0.53 0.37
RoBERTa-large 0.55 0.36 0.61 0.41 0.62 0.43
BART-large 0.47 0.28 0.56 0.35 0.52 0.34
Multi-modal models
Model Acc mAP Acc mAP Acc mAP
CLIP-VIT (text-only) 0.47 0.34 - - - -
CLIP-VIT (retrieval) 0.56 0.35 - - - -
CLIP-VIT (generation) 0.61 0.4 - - - -
ViLT (retrieval) 0.41 0.31 - - - -
ViLT (generation) 0.44 0.32 - - - -
Table 3: Zero-shot performance for affordance prediction using encoder based models. Acc: Accuracy, LM: Language model, VI: Vision. Only LMs are ensembled with VI. The best results are marked in bold.
@@
Table 4: Zero-shot and in-context learning (ICL) performance for affordance prediction using generative models. IR: Image Retrieval; IG: Image Generation. Number of demonstration examples used for ICL = 5. We also mention the variance over different selections of examples. The best results are marked in bold.
5.1 Zero-shot performance
Table 3 shows the results of the zero-shot affordance predictions from the mentioned models. The second column (i.e., Actual) indicates the values from the original LM and multi-modal models. The third and fourth columns (i.e., LM + VI) indicate the performances of ensembling language models with two of the multi-modal models we used. We observe that, the PTLMs have some knowledge about object affordances, but they still lack the comprehensive reasoning ability about these affordances, which is reflected in the low mAP values. Further, the performances vary across different settings. In case of NLI based setup, the fine-tuned RoBERTa and BART models show improvement in the performance, which indicates that during fine-tuning on MNLI dataset, those models gain some reasoning ability. In Table 4 we show the generation based results in a zero-shot setting. In case of FLAN-T5-large model, where we use it to predict a binary label (Yes/No) for an affordance class, the performance drops significantly (the accuracy is less than 7%). This shows that there are still some challenges for the text-to-text models in general reasoning ability about the object affordances. In addition, we find that, the multi-modal models do not perform well in text-only settings, despite being pretrained on text and image data. The performances of the language models get boosted when ensembling with the multi-modal models, which indicates that the prediction of object affordance from sentence is a difficult task, and can be enhanced in presence of images.
5.2 Effect of finetuning on commonsense datasets
We observe that the fine-tuned model on common sense reasoning task (Table 5) show improved performance for the affordance prediction task. This indicates that the pre-trained models lack the reasoning of object affordance. We find that the smallest BERT-base model fine-tuned on PiQA, performs almost similar to that of the BERT-large or BART-large models (see Table 3).
@@ 5.2 Effect of finetuning on commonsense datasets
5.3 Few-shot performance
We find that, in presence of few examples from our affordance dataset, the reasoning capability about object affordances can be enhanced for the PTLMs. The results with 124 shots (62 pairs as discussed earlier) are noted in Table 6. In Table 4, we note the results for the in-context learning performance of the generative LLMs and VLMs. We observe a significant performance gain over zero-shot settings. Having said that, we also observe that, even with the in-context learning, the performance of the generative models (with more than 7b parameters) do not reach even close to the performance of the fine-tuned BERT-large model (340M parameters). This suggests that, for the specific affordance prediction tasks from text, finetuning is absolutely essential even for the state-of-the-art LLMs and VLMs.
NLI based (Normal)
Model Accuracy mAP
RoBERTa-large 0.72 0.49
BART-large 0.69 0.48
MLM based
Model Accuracy mAP
BERT-large-uncased 0.58 0.33
RoBERTa-large 0.77 0.39
BART-large 0.65 0.38
Table 6: Few-shot fine-tuning performances of the PTLMs. Number of training data points used: 124.
6 Error analysis
Encoder based models.
We conducted a qualitative analysis of the erroneous cases for the two models (BART-large and RoBerta-large) in MNLI settings to understand what are the typical causes of errors. We take examples where accuracy is below 0.3. Consider the representative example below.
Sentence: The salt from La Mata is often used as table salt. Object: table salt Top 5 predicted affordances (according to the probability score) - [‘sitOn’, ‘pourFrom’, ‘grasp’, ‘fix’, ‘lookThrough’]
@@ 6 Error analysis
Analysis of generative models.
In Appendix Figure 7(a), we plot the correlation between error rate made by chatGPT for each affordance classes and the classwise annotator agreement. We observe a moderately negative correlation ($\rho=-0.29$) which suggests that there is a chance that the model is making higher mispredictions where the agreement is low. Similarly we observe that the mispredictions made by chatGPT for the most frequent objects has a moderately negative correlation ($\rho=-0.58$) with the annotator agreement. The correlation is shown in Figure 7(b). The trends are similar for the other LLMs. These results together indicate that the those objects and affordance classes which are hard to disambiguate by humans also pose a challenge to the most sophisticated GenAI models in predicting the correct answer.
@@ 7 Conclusion
In this paper we introduced a novel text-based affordance dataset GrAFFORD to investigate the affordance knowledge of pre-trained language models and pre-trained vision language models in different zero-shot settings. Our findings suggest that, the state-of-the-art language models, particularly text-to-text models, still exhibit limitations in their ability to reason about object affordances. Finetuning emerges as the only way to improve in such a complex task and here GrAFFORD promises to be a very valuable resource for researchers. By leveraging our dataset, future studies can contribute to enhancing the reasoning capabilities of language models and advancing the understanding of how language is grounded in the context of objects and their affordances.
@@
We used the publicly available XNLI corpus to curate our GrAFFORD dataset. Our dataset does not contain any harmful or offensive contents. Any personal or sensitive information is anonymized and treated with utmost confidentiality. We ensure the protection of participants’ privacy and obtain informed consent for data collection, annotation, and analysis. We incentivized all the annotators uniformly throughout the annotation process.

### v2

@@ 5 Experiments
We explore various state-of-the-art baselines using pre-trained language models (RoBERTa-large, BART-large), instruction-fine-tuned large language models (e.g., FLAN-T5, Falcon, ChatGPT, Llama-3), pre-trained multi-modal vision and language architectures (CLIP-ViT, ViLT, InstructBLIP, IDEFICS, LLaVA). We observe whether these models gain the knowledge of affordances through their pre-training, fine-tuning on commonsense tasks (NLI, PIQA), or few-shot fine-tuning scenarios.
@@ 5.1.1 Pre-trained language models
We frame the zero-shot prediction task in different ways.
MLM based approach. Here, we pose the zero-shot task as masked word prediction problem. We choose BERT-large-uncased, RoBERTa-large Zhuang et al. (2021), and BART-large Lewis et al. (2020) models for the experiment. We pass the sentence and prompt separated by a [SEP] token as an input to the model. We use the prompt “<Object> can be used for <MASK_TOKEN> by human” and obtain the probability of each affordance label using the logit corresponding to the <MASK_TOKEN>.
[…]
@@ 5.1.2 Commonsense reasoning tasks
To understand whether the injection of the common sense knowledge in the pre-trained models can enhance the performance of the affordance prediction, we first fine-tune the pre-trained models on common sense reasoning dataset such as PIQA Bisk et al. (2019). Then we run the fine-tuned models on our dataset using the MLM setup. We use BERT-base, BERT-large, RoBERTa-large, and BART-large models.
Apart from this, we leverage RoBERTa-large and BART-large fine-tuned on the Multi-genre NLI (MNLI) corpus Williams et al. (2018) to evaluate on NLI setup. We utilize the sentence as premise and use the hypothesis as “<object> can be used for <affordance> by human” for each object-affordance pair, and use the entailment score to rank the affordance classes and report mAP and accuracy. Details of the experiment can be found in Appendix B.1.1.
5.1.3 Multimodal models
@@
Text-only MLM setup
@@
Images contain necessary information about shape, texture, and size of objects that can be utilized to effectively predict an object affordance (such as the handle of the bucket can be used to grasp and lift). Hence, we also convert the problem into a multi-modal task by synthesizing corresponding images from the context sentence, and predict the affordance of an object (mentioned in the sentence) based on the input.
Synthesizing images. In this setup, we use two different techniques to synthesize semantically close images to corresponding context sentences using 1) retrieval and 2) generation. We further use top five images for both, to get an accurate estimation.
Image retrieval: We use the CLIP Radford et al. (2021) based sentence-transformers architecture to search for top five semantically similar images for each of the contexts from the Visualgenome Krishna et al. (2017) dataset.
Image generation: We adopt the generative StableDiffusion Rombach et al. (2022) model to generate top five images based on the sentence as a text prompt. Details can be found in the Appendix B.3.1.
We use the top five retrieved images by using retrieval and generation methods each. We use CLIP Radford et al. (2021) and ViLT Kim et al. (2021) as our vision-text models. CLIP has a text encoder $f_{T}$ and a visual encoder $f_{V}$, which can project text and image into the shared latent space. We aggregate the $k$ (=5) corresponding images and use CLIP to compute the relevance score of (x, y): $Score_{VI}(x,y)=\frac{1}{k}\sum_{i=1}^{K}{\cos{(f_{T}(x),f_{v}(I^{k}_{y}))}}$, where $I^{k}_{y}$ is the $k^{\textrm{th}}$ image for the input text $y$. In the ViLT model, we provide the text prompt along with the representative images as input to predict the masked token. We use the same prompt as the previous MLM task (i.e., “<Object> can be used for <MASK_TOKEN> by human.”) and get the probability of each affordance class as the logit corresponding to the <MASK_TOKEN>.
[…]
Ensemble language and vision prediction.
Following Yang et al. (2022), we use the weighted sum as the late fusion over the final output probabilities of each affordance class from the language and multi-modal models. Experimental details can be found in Appendix B.2.
@@ 5.2 Few-shot prediction
Training data To create few-shot training examples for fine-tuning encoder based PTLMs, we take all the 62 objects, and for each object we randomly select exactly 1 positive affordance class (i.e., the class label annotated as 1) and 1 negative affordance class (i.e., the class label annotated as 0) for generating the training prompt. Overall they constitute 124 training examples (62 sentence-object pairs and 2 selected classes for each) for the few-shot experiment. For more details of the training data curation and the selection of examples for in-context learning, refer to Appendix B.3
@@
Fine-tuning PTLM: We fine-tune the encoder based PTLMs in NLI based setup having the context sentence as premise and use same hypothesis (i.e., “<object> can be used for <affordance> by human”) which we use in the zero-shot settings. We use BERT-large-uncased, RoBERTa-large and BART-large for fine-tuning in this setup. For implementation details refer to Appendix B.3
In-context learning for generative models: We employ the same generative LLMs as well as VLMs to perform affordance prediction using five demonstration examples from the training data. We use the same text prompt as zero-shot setting and concatenate the five demonstration examples along with corresponding label (i.e., ‘NO’ for positive class, and ‘NO’ for the negative class) to the prompt and ask the LLMs and VLMs to predict the affordance. In case of the VLMs, we do not provide any additional image example here.
6 Benchmarking Text2Afford prediction
@@
For each positive class $\in\{P_{1},P_{2},..P_{n}\}$, we compare the predicted scores of that affordance class with the predicted scores of the negative classes $\in\{N_{1},N_{2},..N_{m}\}$. If the predicted score of the positive class is higher than the predicted score of all the negative classes (i.e., $p(P_{i})>p(N_{j})_{\forall j}$), we increment the correct count by 1555During calculation we discard the cases when there is no positive class for a sentence-object pair in the ground truth. We do not find any instance where no negative class is present.. Conversely, if the predicted score of the negative class is higher, we increment the wrong count by 1. The final accuracy is calculated by dividing the total number of correct counts by the total number of the instances. To rank the affordance classes based on the predicted score, we also report the Mean Average Precision (mAP@$K$, where $K$ is the number of affordance classes).
@@
NLI based
Model Actual Fine-tuned LM + VI (CLIP) LM + VI (ViLT)
Acc mAP Acc mAP Acc mAP Acc mAP
RoBERTa-large-mnli 0.64 0.43 0.72 0.49 0.79 0.52 0.79 0.54
BART-large-mnli 0.65 0.38 0.69 0.48 0.62 0.4 0.64 0.43
@@
BERT-large-uncased 0.46 0.26 0.58 0.33 0.55 0.38 0.53 0.37
RoBERTa-large 0.55 0.36 0.77 0.49 0.61 0.41 0.62 0.43
BART-large 0.47 0.28 0.65 0.38 0.56 0.35 0.52 0.34
Multi-modal models (zero-shot)
CLIP-VIT (text-only) 0.47 0.34 - - - - - -
CLIP-VIT (retrieval) 0.56 0.35 - - - - - -
CLIP-VIT (generation) 0.61 0.4 - - - - - -
ViLT (retrieval) 0.41 0.31 - - - - - -
ViLT (generation) 0.44 0.32 - - - - - -
Table 5: Performance for affordance prediction using encoder based models. Acc: Accuracy, LM: Language model, VI: Vision. Only LMs are ensembled with VI. The best results are in bold.
@@
Table 6: Zero-shot and in-context learning (ICL) performance for affordance prediction using generative models. IR: Image Retrieval; IG: Image Generation. Number of demonstration examples used for ICL = 5. We also mention the variance over different selections of examples. The best results are in bold.
[…]
Finetuning on commonsense datasets. We observe that the fine-tuned model on commonsense reasoning task (Table 7) show improved performance for the affordance prediction task. This indicates that the pre-trained models lack the reasoning of object affordance. Interestingly, we find that the smallest BERT-base model fine-tuned on PIQA, performs almost similar to that of the BERT-large or BART-large models (see Table 5).
@@
Few-shot performance. We find that, in presence of few examples from our affordance dataset, the reasoning capability about object affordances can be enhanced for the PTLMs. The results with 124 shots (62 pairs as discussed earlier) are noted in Table 5. In Table 6, we note the results for the in-context learning performance of the generative LLMs and VLMs. We observe a significant performance gain over zero-shot settings. Having said that, we also observe that, even with the in-context learning, the performance of the generative models (with more than 7b parameters) do not reach even close to the performance of the fine-tuned BERT-large model (340M parameters). This suggests that, for the specific affordance prediction tasks from text, finetuning is absolutely essential even for the state-of-the-art LLMs and VLMs.
Error analysis
Encoder based models. We conducted a qualitative analysis of the erroneous cases for the two models (BART-large and RoBerta-large) in MNLI settings to understand what are the typical causes of errors. We take examples where accuracy is below 0.3. Consider the representative example below.
{mdframed}
[backgroundcolor=gray!10] Sentence: The salt from La Mata is often used as table salt. Object: table salt
Top 5 predicted affordances (according to the probability score) - [‘sitOn’, ‘pourFrom’, ‘grasp’, ‘fix’, ‘lookThrough’]
@@
Analysis of generative models.  In Appendix Figure 6(a), we plot the correlation between error rate made by chatGPT for each affordance classes and the classwise annotator agreement. We observe a moderately negative correlation ($\rho=-0.29$) which suggests that there is a chance that the model is making higher mispredictions where the agreement is low. Similarly we observe that the mispredictions made by chatGPT for the most frequent objects has a moderately negative correlation ($\rho=-0.58$) with the annotator agreement. The correlation is shown in Figure 6(b). The trends are similar for the other LLMs. These results together indicate that those objects and affordance classes which are hard to disambiguate by humans also pose a challenge to the most sophisticated GenAI models in predicting the correct answer.
7 Text2Afford for physical reasoning
Apart from benchmarking LLMs and VLMs, we observe whether Text2Afford can be used as a source of affordance knowledge. We choose the physical commonsense reasoning as a target as the ‘Object affordance’ represents an innate physical property of an object, and we believe that any language model having strong affordance reasoning capability can enhance the physical reasoning capability. To explore this, we perform an ‘instruction fine-tuning’ on the Text2Afford dataset (although it is not meant for training) using few open-source LLMs (llama-3-8b-instruct, flan-t5), and test on two physical reasoning dataset - (1) PROST Aroca-Ouellette et al. (2021), which contains 10 types of different physical properties of an object (including 6 affordance properties - rolling, breaking, stacking, grasping, sliding, bouncing) along with complex reasoning questions, and (2) PIQA Bisk et al. (2019) which, focuses on selecting appropriate option given a situation that requires physical commonsense.
For PROST, using llama-3 the accuracy boosts from 0.36 to 0.42 after instruct fine-tuning with Text2Afford. Moreover, out of the 6 affordance properties from PROST, the accuracy got boosted for the reasoning of 5 affordance properties. For the PIQA, the same LLM gives a maximum of 4% accuracy boost. The full result is shown in Table 8. This suggest the generalizability of Text2Afford in physical reasoning tasks.
Model Dataset
PROST PIQA
Zero-shot +Text2Afford Zero-shot +Text2Afford
Llama-3-8b 0.36 0.42(+.06) 0.74 0.78(+.04)
FLAN-T5-xl 0.13 0.16(+.03) 0.57 0.59(+.02)
FLAN-T5-xxl 0.34 0.38(+.04) 0.72 0.75(+.03)
Table 8: Text-only physical reasoning dataset evaluation using different LLMs fine-tuned on Text2Afford. +Text2Afford: instruction fine-tuned on Text2Afford
@@ 8 Conclusion
In this paper we introduced a novel text-based affordance dataset Text2Afford to investigate the affordance knowledge of PTLMs and pre-trained VLMs in different zero-shot settings. Our findings suggest that, the state-of-the-art language models, particularly text-to-text models, still exhibit limitations in their ability to reason about object affordances. In this seemingly easy task, we observe how context can introduce various levels of ambiguity and difficulty. We also observe, that even in the presence of such difficulty, human performance is superior and LLMs/VLMs still face difficulty in gaining such knowledge during their pretraining. Additionally, we observe how our dataset provides some additional knowledge that can be useful for physical commonsense reasoning – stressing its orthogonality more with respect to the pretraining knowledge LLMs and VLMs possess.
@@
We used the publicly available XNLI corpus to curate our Text2Afford dataset. Our dataset does not contain any harmful or offensive contents. Any personal or sensitive information is anonymized and treated with utmost confidentiality. We ensure the protection of participants’ privacy and obtain informed consent for data collection, annotation, and analysis. We incentivized all the annotators uniformly throughout the annotation process.
Appendices